OpenStudy (anonymous):

If the pattern is continued what will be the next three numbers? 4,4,8,12,20,32 What will be the eleventh number?

OpenStudy (freckles):

4+4=? 4+8=? 8+12=? 12+20=?

OpenStudy (anonymous):

I feel like a idiot right now thanks :)

OpenStudy (freckles):

sorry didn't mean to make you feel like an idiot :p